# MMR-Mamba: Multi-Contrast MRI Reconstruction with Mamba and Spatial-Frequency Information Fusion

### v1

@@ III-C Target-guided Cross Mamba
For the extracted target modality feature $F_{tar}$ and auxiliary modality feature $F_{aux}$, complementary information is distributed across different regions, making efficient integration crucial for reconstructing high-quality images from under-sampled data. Existing fusion methods based on CNNs and Transformers face challenges due to their restricted capability in handling long-range dependencies or their high computational requirements. To address this, we leverage state space models for target and auxiliary feature fusion, benefiting from their ability to model long-range dependencies. Inspired by the cross model Mamba in [46], we design the Target-guided Cross Mamba (TCM) module to enrich the target modality features with complementary information from auxiliary features.
Specifically, we derive $z_{tar}$ and $z_{aux}$ from the target modality feature $F_{tar}$ and auxiliary modality feature $F_{aux}$ using normalization layer and linear layer:
@@
$z_{aux}=Linear(Norm(F_{aux})).$
Then $z_{tar}$ and $z_{aux}$ are projected into the hidden state space through one-dimensional convolution with the SiLU activation function and SSM without gating as
@@
$H_{aux}=SSM(SiLU(Conv1d(z_{aux}))).$
After that, to maximally restore the target modality information and selectively integrate the auxiliary information, we adopt $z_{tar}$ as the gating to modulate hidden state features $H_{tar}$ and $H_{aux}$, and implement the hidden state feature fusion as
$F_{spa}=H_{tar}\otimes SiLU(z_{tar})+H_{aux}\otimes SiLU(z_{tar}),$ (13)
where $\otimes$ represents element-wise production. In this way, we build the TCM module in a hidden state space based on the gating mechanism. The gating parameters are derived from the target modality, ensuring that the target modality predominates in the fusion process and selectively incorporates complementary information from the auxiliary modality.
@@ III-D Selective Frequency Fusion
To further fuse the complementary information from the extracted features in a broader view, we resort to the frequency domain (i.e.,  Fourier domain), where each frequency component corresponds to all the pixels in the spatial domain, inherently capturing global properties. Additionally, as highlighted in [44, 45], the phase spectrum in the Fourier domain preserves global structure well, while the image degradation mainly manifests in the amplitude spectrum. To comprehensively integrate the complementary information and to restore the degraded features in the amplitude spectrum, we propose Selective Frequency Fusion (SFF) module within the frequency domain, illustrated in Fig. 2.
Given the extracted features $F_{tar}$ and $F_{aux}$ from previous modules, we first transform them to frequency domain through Fourier transform and obtain their amplitude spectrum and phase spectrum:
@@
$\mathcal{A}_{aux},\mathcal{P}_{aux}$ $=\mathcal{F}(F_{aux}).$
For the phase spectrum fusion, we perform element-wise addition on $\mathcal{P}_{aux}$ and $\mathcal{P}_{tar}$, as both contain crucial and consistent structure information [47, 48]. Regarding the amplitude spectrum, it encapsulates style information, which varies significantly across different modalities. Moreover, the amplitude of the under-sampled low-quality images contains interference information that can negatively impact the final reconstruction. Direct concatenation of the amplitude spectrum from these modalities would introduce incompatible and interference information, compromising the quality of the reconstructed image [47, 48].
Figure 2: Illustration of Selective Frequency Fusion (SFF) module.
To mitigate this issue, we adopt a selective strategy to fuse the amplitude spectrum. Specifically, we design a Selective Amplitude Harmonization Module (SAHM), wherein the amplitude spectrum across modalities is dynamically adjusted based on global statistics. At first, we conduct element-wise addition of $\mathcal{A}_{aux}$ and $\mathcal{A}_{tar}$ to produce an intermediate feature $\mathcal{A}$. Next, we adopt global average pooling to generate channel-wise statistics $s\in\mathbb{R}^{C}$, embedding global information. Specifically, the $c-th$ element of $s$ is calculated by shrinking $\mathcal{A}$ through spatial dimensions $H\times W$. For computational efficiency, a fully connected (FC) layer is utilized to create a compact feature $z\in\mathbb{R}^{L\times 1}$, which is further used to guide the adaptive selection. The operation can be expressed as follows:
@@
where $A,B\in\mathcal{R}^{C\times L}$ and $a,~{}b$ represent the soft attention vector for $\mathcal{A}_{aux}$ and $\mathcal{A}_{tar}$. At last, the fused amplitude spectrum $\overline{\mathcal{A}}$ is obtained by weighted sum of $\mathcal{A}_{aux}$ and $\mathcal{A}_{tar}$:
$\overline{\mathcal{A}_{c}}=a_{c}\cdot\mathcal{A}_{aux}+b_{c}\cdot\mathcal{A}_{tar}.$ (17)
@@
$\mathcal{P}^{\prime}=ConvP(\mathcal{P}_{aux}+\mathcal{P}_{tar}).$
@@
where $\mathcal{F}^{-1}$ donates inverse Fourier transform. Through the FFT and selective fusion in the Fourier domain, the SFF module efficiently achieves global feature fusion and recovers high-frequency signals for structure details.
Figure 3: Illustration of the Adaptative Spatial-Frequency Fusion module.
III-E Adaptative Spatial-Frequency Fusion
To further enhance the fused spatial domain feature $F_{spa}$ and fused frequency domain feature $F_{fre}$, and to facilitate the efficient integration of these features, we introduce the Adaptative Spatial-Frequency Fusion (ASFF) module, as illustrated in Fig. 3. The ASFF module allows the two domains to mutually complement each other through channel-wise adaptive integration according to the informativeness of the channel, which is measured by the Batch-Normalization (BN) layer. Previous studies [49, 50] have shown that the BN value indicates channel importance, with lower norms signifying lesser relevance to the final outcomes. Consequently, we propose that channels with diminished significance in one domain can be reinforced by incorporating corresponding channel information from the other domain.
Firstly, we measure channel-wise informativeness of the fused spatial domain feature $F_{spa}$ and frequency feature $F_{fre}$ through the BN layer. Denote the $c-th$ channel of the features as $F_{spa,c}$ and $F_{fre,c}$, the BN is computed as follows:
@@ III-E Adaptative Spatial-Frequency Fusion
where $\omega_{spa,c}$ and $\omega_{fre,c}$ are trainable scaling factors; $\beta_{spa,c}$ and $\beta_{fre,c}$ are trainable offset, and $\epsilon$ is a small constant to avoid divisions by zero.
@@
where $\otimes$ donates element-wise multiplication and $\tau_{spa}$ and $\tau_{fre}$ are the thresholds, we set it according to the maximum and the minimum of the scaling factor as follows:
@@
where $\alpha$ is a value, we set it as 0.1 in our experiments.
Through the ASFF module, less informative features from one domain are enhanced by the corresponding features from the other, while redundant information is effectively eliminated. This process ensures that each channel contributes optimally to the final fused features. By leveraging the most relevant information from spatial and frequency domains, the ASFF module produces feature representations that are most conducive to generating high-quality target images. After the channel-wise incorporation, the features $F_{spa}^{\prime}$ and $F_{fre}^{\prime}$ are concatenated and fed to CNN Decoder to recover the reconstructed target modality image $\hat{I}_{tar}$:
@@ IV-A Dataset Description
In the study, two datasets with different anatomical structures and protocols are adopted for evaluation, including the BraTS dataset [51] and fastMRI knee dataset [52]. The BraTS Dataset contains both T1WIs and T2WIs scans of the brain. We extract 2D images uniformly from 100 3D MRI volumes in the BraTS dataset. The dataset is split subject-wise with a 3:1 ratio, yielding a total of 3,621 images for training and 1,088 images for testing. The 2D image size is $240\times 240$. In our experiments, we adopt T1WI as the auxiliary modality for the reconstruction of the T2WI modality. The fastMRI Dataset is the largest public MRI dataset with raw k-space data. Following [53], 227 and 45 pairs of single-coil PDWI and FS-PDWI knee volumes are selected for training and testing, respectively, resulting in a total of 8,332 pairs of 2D images for training and 1,665 images for testing. The 2D image size is $320\times 320$. In our experiments, we adopt PDWI as the auxiliary modality for the reconstruction of the FS-PDWI modality.
@@ IV-B Experimental Setup
Comparison Methods. To demonstrate the effectiveness of our proposed MMR-Mamba, we compare it against seven multi-contrast reconstruction methods: MDUNet [5], MINet [54], MCCA [13], MTrans [4], DCAMSR [14], SwinIR [55], and Pan-Mamba [46]. For a fair comparison, we run all the models for $100,000$ iterations with a batch size of 4. Other settings of the compared methods followed the implementations provided in their original papers.
Performance Metrics. For quantitative evaluation, we assess the image reconstruction results using peak-to-noise ratio (PSNR), structural similarity index (SSIM), and normalized mean squared error (NMSE). For qualitative evaluation, we visualize the reconstruction results with error maps.
Implementation Details. The proposed framework is developed with PyTorch, and the training and testing processes are executed on one NVIDIA RTX A6000 GPU (48GB). We employ an AdamW optimizer with an initial learning rate of $0.001$ and a weight decay of $41e-$ 4 for the training. We run $100,000$ iterations with a batch size of 4 for both datasets. For both datasets, the undersampled images are obtained by applying a 1D cartesian random under-sample mask with 4×4$\times$4 × and 8×8$\times$8 × acceleration.
TABLE I: Quantitative results on the BraTS and fastMRI datasets with different acceleration factors. We report mean±std for the PSNR, SSIM, and NMSE metrics. The best results are highlighted in red.
BraTS 4×4$\times$4 × 8×8$\times$8 ×
@@
Zero-filling [56] JMRI’01 30.11±1.53 0.767±0.038 5.659±0.018 26.58±1.49 0.673±0.036 12.469±0.033
MDUNet [5] TBME’18 37.94±1.66 0.975±0.006 0.905±0.003 35.19±1.64 0.960±0.009 1.690±0.005
MINet [54] MICCAI’21 38.26±1.74 0.976±0.006 0.847±0.003 35.23±1.72 0.961±0.009 1.697±0.005
SwinIR [55] CVPR’21 37.87±1.73 0.974±0.006 0.926±0.003 34.95±1.72 0.960±0.009 1.805±0.005
MTrans [4] TMI’22 36.02±1.67 0.962±0.007 1.429±0.005 34.81±1.57 0.957±0.009 1.868±0.005
MCCA [13] JBHI’23 38.03±1.68 0.975±0.006 0.903±0.003 35.37±1.66 0.962±0.009 1.633±0.005
DCAMSR [14] MICCAI’23 38.60±1.75 0.978±0.006 1.193±0.004 35.99±1.74 0.965±0.009 1.417±0.005
Pan-Mamba [46] Arxiv’24 38.84±1.79 0.978±0.006 0.739±0.273 36.18±1.77 0.966±0.009 1.363±0.466
MMR-Mamba (Ours) 40.98±1.88 0.985±0.005 0.454±0.190 37.75±1.85 0.974±0.008 0.955±0.358
fastMRI 4×4$\times$4 × 8×8$\times$8 ×
@@
Zero-filling [56] JMRI’01 27.68±1.75 0.571±0.061 5.020±0.013 25.64±1.64 0.454±0.070 7.719±0.02
MDUNet [5] TBME’18 28.60±1.00 0.600±0.050 4.000±0.010 27.90±0.86 0.544±0.050 4.600±0.02
MINet [54] MICCAI’21 29.47±1.88 0.639±0.069 3.334±0.009 28.17±1.74 0.563±0.081 4.327±0.01
SwinIR [55] CVPR’21 29.42±1.87 0.636±0.069 3.382±0.009 28.09±1.74 0.560±0.081 4.407±0.01
MTrans [4] TMI’22 29.00±1.79 0.619±0.068 3.699±0.009 27.31±1.68 0.526±0.081 5.254±0.01
MCCA [13] JBHI’23 29.46±1.87 0.637±0.069 3.346±0.009 28.23±1.75 0.562±0.081 4.275±0.01
DCAMSR [14] MICCAI’23 29.45±1.87 0.637±0.068 3.349±0.009 28.42±1.79 0.569±0.081 4.098±0.01
Pan-Mamba [46] Arxiv’24 29.59±1.87 0.645±0.068 3.248±0.928 28.36±1.76 0.570±0.081 4.159±1.034
MMR-Mamba (Ours) 29.66±1.88 0.647±0.068 3.201±0.927 28.44±1.76 0.572±0.081 4.084±1.036
@@ IV-C Experimental Results
Quantitative Results. In Table I, we report the PSNR, SSIM, and NMSE results on both datasets for 4$\times$ and 8$\times$ acceleration. Firstly, focusing on the BraTS dataset in the upper part of the table, our method achieves the best reconstruction results, achieving a PSNR of 40.98 dB and SSIM of 0.985 under 4$\times$ acceleration, and a PSNR of 37.75 dB and SSIM of 0.974 under 8$\times$ acceleration. These results highlight our model’s efficacy in fusing information from multi-contrast images for reconstructing target images. Additionally, our method outperforms the second-best method, Pan-Manba, by 2.14 dB in PSNR under 4$\times$ acceleration and by 1.57 dB in PSNR under 8$\times$ acceleration. The fact that Pan-Manba achieves the second-best results further highlights the effectiveness of the Mamba in MRI reconstruction. Similarly, as shown in the lower part of the table, our method achieves the best reconstruction results on the fastMRI knee dataset, with a PSNR of 29.66 dB and SSIM of 0.647 under 4$\times$ acceleration and a PSNR of 28.44 dB and SSIM of 0.572 under 8$\times$ acceleration. This further demonstrates the effectiveness and robustness of our model in reconstructing high-quality MR images.

### v2

@@ III-C Target-guided Cross Mamba
As correlated features distribute across different regions and modalities, capturing and fusing these correlated features is crucial for the reconstruction of high-quality target modality images. Existing fusion methods based on CNNs and Transformers face challenges due to their restricted capability in handling long-range dependencies or their high computational requirements. To address this, we leverage state space models for target and reference feature fusion, benefiting from their ability to efficiently model long-range dependencies. Inspired by the cross-modal Mamba in [49], we design the Target-guided Cross Mamba (TCM) module to enrich the target modality features with complementary information from reference features.
Specifically, we derive $z_{tar}$ and $z_{ref}$ from the target modality feature $F_{tar}$ and reference modality feature $F_{ref}$ using a normalization layer and a linear layer:
@@
$z_{ref}=Linear(Norm(F_{ref})).$
Then $z_{tar}$ and $z_{ref}$ are projected into the hidden state space through one-dimensional convolution with the SiLU activation function and SSM without gating as
@@
$H_{ref}=SSM(SiLU(Conv1d(z_{ref}))).$
After that, to maximally restore the target modality information and selectively integrate the reference information, we adopt $z_{tar}$ as the gating to modulate hidden state features $H_{tar}$ and $H_{ref}$, and implement the hidden state feature fusion as
$F_{spa}=H_{tar}\otimes SiLU(z_{tar})+H_{ref}\otimes SiLU(z_{tar}),$ (13)
where $\otimes$ represents element-wise production. In this way, we build the TCM module in a hidden state space based on the gating mechanism. The gating parameters are derived from the target modality, ensuring that the target modality predominates in the fusion process and selectively incorporates complementary information from the reference modality.
@@ III-D Selective Frequency Fusion
To further fuse the complementary information from the extracted features in a broader view, we resort to the frequency domain (i.e.,  Fourier domain), where each frequency component corresponds to all the pixels in the spatial domain, inherently capturing global properties. Additionally, as highlighted in [47, 48], the phase spectrum in the Fourier domain preserves global structure well, while the image degradation mainly manifests in the amplitude spectrum. To comprehensively integrate the complementary information and to restore the degraded features in the amplitude spectrum, we propose the Selective Frequency Fusion (SFF) module within the frequency domain, illustrated in Fig. 2.
Given the extracted features $F_{tar}$ and $F_{ref}$ from previous modules, we first transform them to frequency domain through Fourier transform and obtain their amplitude spectrum and phase spectrum:
@@
$\mathcal{A}_{ref},\mathcal{P}_{ref}$ $=\mathcal{F}(F_{ref}).$
For the phase spectrum fusion, we perform element-wise addition on $\mathcal{P}_{ref}$ and $\mathcal{P}_{tar}$, as both contain crucial and consistent structural information [50, 51]. Regarding the amplitude spectrum, it encapsulates style information, which varies significantly across different modalities. Moreover, the amplitude of the under-sampled low-quality images contains interference information that can negatively impact the final reconstruction. Direct concatenation of the amplitude spectrum from these modalities would introduce incompatible and interference information, compromising the quality of the reconstructed image [50, 51].
To mitigate this issue, we adopt a selective strategy to fuse the amplitude spectrum. Specifically, we design a Selective Amplitude Harmonization Module (SAHM), wherein the amplitude spectrum across modalities is dynamically adjusted based on global statistics. First, we conduct element-wise addition of $\mathcal{A}_{ref}$ and $\mathcal{A}_{tar}$ to produce an intermediate feature $\mathcal{A}$. Next, we adopt global average pooling to generate channel-wise statistics $s\in\mathbb{R}^{C}$, embedding global information. Specifically, the $c-th$ element of $s$ is calculated by shrinking $\mathcal{A}$ through spatial dimensions $H\times W$. For computational efficiency, a fully connected (FC) layer is utilized to create a compact feature $z\in\mathbb{R}^{L\times 1}$, which is further used to guide the adaptive selection. The operation can be expressed as follows:
@@
where $A,B\in\mathcal{R}^{C\times L}$ and $a,~{}b$ represent the soft attention vector for $\mathcal{A}_{ref}$ and $\mathcal{A}_{tar}$. Then, the fused amplitude spectrum $\overline{\mathcal{A}}$ is obtained by weighted sum of $\mathcal{A}_{ref}$ and $\mathcal{A}_{tar}$:
$\overline{\mathcal{A}_{c}}=a_{c}\cdot\mathcal{A}_{ref}+b_{c}\cdot\mathcal{A}_{tar}.$ (17)
Figure 2: Illustration of Selective Frequency Fusion (SFF) module.
@@
$\mathcal{P}^{\prime}=ConvP(\mathcal{P}_{ref}+\mathcal{P}_{tar}).$
@@
where $\mathcal{F}^{-1}$ donates inverse Fourier transform (IFT). Through the FT and selective fusion in the Fourier domain, the SFF module efficiently achieves global feature fusion and recovers high-frequency signals for reconstructing structural details.
Figure 3: Illustration of the Adaptive Spatial-Frequency Fusion module.
III-E Adaptive Spatial-Frequency Fusion
To further enhance the fused spatial domain feature $F_{spa}$ and fused frequency domain feature $F_{fre}$, and to facilitate the efficient integration of these features, we introduce the Adaptive Spatial-Frequency Fusion (ASFF) module, as illustrated in Fig. 3. The ASFF module allows the two domains to mutually complement each other through channel-wise adaptive integration based on the informativeness of each channel, which is measured by the Batch-Normalization (BN) layer. Previous studies [52, 53] have shown that the BN value indicates channel importance, with lower norms signifying lesser relevance to the final outcomes. Consequently, we propose that channels with diminished significance in one domain can be reinforced by incorporating corresponding channel information from the other domain.
Firstly, we measure the channel-wise informativeness of the fused spatial domain feature $F_{spa}$ and frequency feature $F_{fre}$ through the BN layer. Denote the $c-th$ channel of the features as $F_{spa,c}$ and $F_{fre,c}$, the BN is computed as follows:
@@ III-E Adaptive Spatial-Frequency Fusion
where $\omega_{spa,c}$ and $\omega_{fre,c}$ are trainable scaling factors; $\beta_{spa,c}$ and $\beta_{fre,c}$ are trainable offsets, and $\epsilon$ is a small constant to avoid division by zero.
@@
where $\otimes$ denotes element-wise multiplication and $\tau_{spa}$ and $\tau_{fre}$ are the thresholds, set according to the maximum and the minimum of the scaling factor as follows:
@@
where $\alpha$ is a value set to 0.1 in our experiments.
Through the ASFF module, less informative features from one domain are enhanced by the corresponding features from the other, while redundant information is effectively eliminated. This process ensures that each channel contributes optimally to the final fused features. By leveraging the most relevant information from both spatial and frequency domains, the ASFF module produces feature representations that are most conducive to generating high-quality target images. After the channel-wise incorporation, the features $F_{spa}^{\prime}$ and $F_{fre}^{\prime}$ are concatenated and fed to the CNN Decoder to recover the reconstructed target modality image $\hat{I}_{tar}$:
@@ IV-A Dataset Description
In this study, we adopt two datasets with different anatomical structures and protocols for evaluation: the BraTS dataset [3] and the fastMRI knee dataset [54]. The BraTS Dataset contains both T1WIs and T2WIs scans of the brain. We extract 2D images uniformly from 100 3D MRI volumes in the BraTS dataset. The dataset is split subject-wise with a 3:1 ratio, yielding a total of 3,621 images for training and 1,088 images for testing. The 2D image size is $240\times 240$. In our experiments, we use T1WI as the reference modality for the reconstruction of the T2WI modality. The fastMRI Dataset is the largest public MRI dataset with raw k-space data. Following [55], we select 227 and 45 pairs of single-coil PDWI and FS-PDWI knee volumes for training and testing, respectively, resulting in a total of 8,332 pairs of 2D images for training and 1,665 images for testing. The 2D image size is $320\times 320$. In our experiments, we use PDWI as the reference modality for the reconstruction of the FS-PDWI modality.
@@ IV-B Experimental Setup
Comparison Methods. To demonstrate the effectiveness of our proposed MMR-Mamba, we compare it against seven multi-modal reconstruction methods: MDUNet [10], MINet [56], MCCA [16], MTrans [2], DCAMSR [17], SwinIR [57], and Pan-Mamba [49]. For a fair comparison, we run all the models for $100,000$ iterations with a batch size of 4. Other settings of the compared methods followed the implementations provided in their original papers.
Performance Metrics. For quantitative evaluation, we assess the image reconstruction results using peak signal-to-noise ratio (PSNR), structural similarity index (SSIM), and normalized mean squared error (NMSE). For qualitative evaluation, we visualize the reconstruction results with error maps.
Implementation Details. The proposed framework is developed with PyTorch, and the training and testing processes are executed on an NVIDIA RTX A6000 GPU (48GB). We employ an AdamW optimizer with an initial learning rate of $0.001$ and a weight decay of $41e-$ 4 for the training. We run $100,000$ iterations with a batch size of 4 for both datasets. For both datasets, the undersampled images are obtained by applying a 1D Cartesian random under-sample mask with 4×4$\times$4 × and 8×8$\times$8 × acceleration.
TABLE I: Quantitative results on the BraTS and fastMRI datasets with different acceleration factors. We report mean±std for the PSNR, SSIM, and NMSE metrics, along with network parameters. The best results are highlighted in red.
BraTS 4×4$\times$4 × 8×8$\times$8 × Param (M)
@@
Zero-filling [58] JMRI’01 30.11±1.53 0.767±0.038 5.659±1.816 26.58±1.49 0.673±0.036 12.469±3.337 -
MDUNet [10] TBME’18 37.94±1.66 0.975±0.006 0.905±0.310 35.19±1.64 0.960±0.009 1.690±0.519 7.40
MINet [56] MICCAI’21 38.26±1.74 0.976±0.006 0.847±0.312 35.23±1.72 0.961±0.009 1.697±0.568 3.75
SwinIR [57] CVPR’21 37.87±1.73 0.974±0.006 0.926±0.334 34.95±1.72 0.960±0.009 1.805±0.590 2.99
MTrans [2] TMI’22 36.02±1.67 0.962±0.007 1.429±0.496 34.81±1.57 0.957±0.009 1.868±0.574 746.02
MCCA [16] JBHI’23 38.03±1.68 0.975±0.006 0.903±0.312 35.37±1.66 0.962±0.009 1.633±0.511 15.21
DCAMSR [17] MICCAI’23 38.60±1.75 0.978±0.006 1.193±0.402 35.99±1.74 0.965±0.009 1.417±0.464 10.83
Pan-Mamba [49] Arxiv’24 38.84±1.79 0.978±0.006 0.739±0.273 36.18±1.77 0.966±0.009 1.363±0.466 2.12
MMR-Mamba (Ours) 40.98±1.88 0.985±0.005 0.454±0.190 37.75±1.85 0.974±0.008 0.955±0.358 0.42
fastMRI 4×4$\times$4 × 8×8$\times$8 × Param (M)
@@
Zero-filling [58] JMRI’01 27.68±1.75 0.571±0.061 5.020±1.271 25.64±1.64 0.454±0.070 7.719±1.594 -
MDUNet [10] TBME’18 28.60±1.00 0.600±0.050 4.000±1.021 27.90±0.86 0.544±0.050 4.600±1.942 7.40
MINet [56] MICCAI’21 29.47±1.88 0.639±0.069 3.334±0.937 28.17±1.74 0.563±0.081 4.327±1.019 3.75
SwinIR [57] CVPR’21 29.42±1.87 0.636±0.069 3.382±0.949 28.09±1.74 0.560±0.081 4.407±1.032 2.99
MTrans [2] TMI’22 29.00±1.79 0.619±0.068 3.699±0.929 27.31±1.68 0.526±0.081 5.254±1.049 746.02
MCCA [16] JBHI’23 29.46±1.87 0.637±0.069 3.346±0.937 28.23±1.75 0.562±0.081 4.275±1.019 15.21
DCAMSR [17] MICCAI’23 29.45±1.87 0.637±0.068 3.349±0.935 28.42±1.79 0.569±0.081 4.098±1.027 10.83
Pan-Mamba [49] Arxiv’24 29.59±1.87 0.645±0.068 3.248±0.928 28.36±1.76 0.570±0.081 4.159±1.034 2.12
MMR-Mamba (Ours) 29.66±1.88 0.647±0.068 3.201±0.927 28.44±1.76 0.572±0.081 4.084±1.036 0.42
@@ IV-C Experimental Results
Quantitative Results. In Table I, we report the PSNR, SSIM, and NMSE results on both datasets for 4$\times$ and 8$\times$ acceleration. Firstly, focusing on the BraTS dataset in the upper part of the table, our method achieves the best reconstruction results, with a PSNR of 40.98 dB and SSIM of 0.985 under 4$\times$ acceleration, and a PSNR of 37.75 dB and SSIM of 0.974 under 8$\times$ acceleration. These results highlight our model’s efficacy in fusing information from multi-modal images for reconstructing target images. Additionally, our method outperforms the second-best method, Pan-Mamba, by 2.14 dB in PSNR under 4$\times$ acceleration and by 1.57 dB in PSNR under 8$\times$ acceleration. The fact that Pan-Mamba achieves the second-best results further highlights the effectiveness of the Mamba in MRI reconstruction. Similarly, as shown in the lower part of the table, our method achieves the best reconstruction results on the fastMRI knee dataset, with a PSNR of 29.66 dB and SSIM of 0.647 under 4$\times$ acceleration and a PSNR of 28.44 dB and SSIM of 0.572 under 8$\times$ acceleration. This further demonstrates the effectiveness and robustness of our model in reconstructing high-quality images.